# Planning Shorter Paths in Graphs of Convex Sets by Undistorting Parametrized Configuration Spaces

### v1

@@ III-D2 Optimal Step Sizes
Our objectives are too complex to easily identify the Lipschitz constant and theoretically determine a good step size. To avoid experimental testing and manual tuning of step size in classical PGD, we use a variant that searches for and uses an optimal step size. In particular, we use backtracking line search PGD [27], which starts with an upper bound on the step size, and repeatedly halves it until the Armijo condition of sufficient decrease is met. This keeps the solver from overshooting minima while maintaining speed in convergence.
@@ III-D3 Gradient Precompilation
Initially, gradient computations were the majority of the runtime. Precompiling gradients with JAX [28] moves this time cost offline, speeding up the PGD iterations quick. By precompiling gradients for each vertex individually, we ensure we can use the same computations for any start and goal.
@@ III-E More general nonconvex objectives: Curvature
should help TOPP produce better trajectories, as high curvature paths contain tight turns, that require a slower traversal to stay within acceleration limits. Although such paths might be longer than those produced by a pure shortest-path trajectory, they can be traversed more quickly.
Similar to the cases focusing on undistorting paths, we define this objective over sampled points along the path defined by $x_{v}$. Given sampled points, we calculate the curvature of each point and then apply the RealSoftMax (a smooth maximum function) to approximate the maximum curvature of our paths. We expect paths under this optimization will have higher path length but lower duratrion when time-parametrized by TOPP.
@@ IV-A Constrained Bimanual Motion Planning
In this experiment two iiwas navigate a shelf while keeping the transform between end effectors constant, as if they were jointly carrying an object. This setup, shown in Figure 1, matches the setup in [3]. We evaluate the PGD solver on the key start and goal pairs from [3]. The comparison of these benchmark paths before and after optimizing for the nonconvex objective is presented in Table I. The “GCS” column indicates using the convex $\mathbb{R}^{8}$ objective, the “Distance” column indicates using the nonconvex $\mathbb{R}^{14}$ objective, and the “Curvature + Distance” column indicates a linear combination of the nonconvex $\mathbb{R}^{14}$ objective and the nonconvex curvature cost with a ratio of 8 to 0.01 respectively. While the $\mathbb{R}^{14}$ objective results in the shortest paths, regularizing for lower curvature lengthens paths, but shortens traversal times after TOPP’s re-timing. Visually, minimizing this joint objective leads to more rounded paths, as shown in Figure 4.
To quantify the difference in distance traveled between the arms, we define the imbalance of a trajectory as $(d_{s}-d_{c})/(d_{s}+d_{c})$, where $d_{c}$ is the distance traveled by the controlled arm and $d_{s}$ is the distance traveled by the subordinate arm. When both arms travel comparable distances, the imbalance distribution centers around 0. When one arm travels much longer distributions than the other, the imbalance metric approaches ±$\pm 1$ in magnitude. As indicated by Table I this imbalance metric approaches 0 after post-processing under the $\mathbb{R}^{14}$ objective. This shift is also apparent in Figure 3 as the path becomes more centered. The imbalance for the joint curvature and distance optimization is higher than the $\mathbb{R}^{14}$ distance optimization indicating that smoother paths are more imbalanced. This asymmetry likely comes from the same-handedness of the iiwas.
TABLE I: Optimizing over the nonconvex cost improves metrics for the three genchmark trajectories.
@@
Imbalance 0.493 0.199 0.217
Figure 3: Paths become more centered because the nonconvex objective used by the PGD considers the distance traveled by both arms, rather than just the controlled arm in the original convex GCS objective.
For a more comprehensive analysis, we randomly sample 100 start and end points from the valid and reachable configuration space. Paths generated are on average 20.60% shorter in the $\mathbb{R}^{14}$ configuration space after applying our postprocessing step. The total time taken to navigate these paths is on average 31.02% shorter. The imbalance shifts towards 0, indicating that the paths for the subordinate arm are more comparable to the leading arm after the nonconvex optimization. These improvements took an average of 0.0554 seconds of compute (approximately 13.7 iterations that took varying time based on length and complexity of the path) in addition to the 2.133 seconds that the surrogate convex optimization takes.
@@
Figure 4: Paths optimized jointly for curvature and distance result in quicker trajectories, even though they are longer (as seen near the top). Note that the curvature-regularized path does not pass as close to the edge of the shelf.
@@ IV-B Planning over $\operatorname{SO}(3)$
To plan shortest paths over $\operatorname{SO}(3)$, we set up the same charts and convex regions as in [4] for the Euler angles, quaternions, and axis-angle parametrizations. The latter two leverage piecewise-linear approximations of the original $\operatorname{SO}(3)$ space and are there for reference. We run PGD on the Euler angles setting only. For Euler angles, the GCS graph is fully connected and is optimizing Euclidean distance within each set, so the shortest path between any two points will be a linear path, regardless of the order of our Bézier curves. For time-efficiency, we generate order one GCS solutions and initialize the PGD solver with the control points evenly spaced along each line segment.
Given there are no obstacles, SLERP gives the shortest path between any two orientations in this setting. We use it for a closed-form ground truth distance for each start and goal pair. Figure 5 shows the distribution of relative error in the path length when using the three representations of $\operatorname{SO}(3)$ to plan across 125 random start and goal pairs, along with the PGD post-processing on the Euler angles paths. The distribution of error for Euler angles shifts significantly closer to 0 after running PGD on it. The relative error decreases by 42.5% on average. This improvement has a bimodal distribution: for some paths the PGD greatly improves the solution, whereas for others, there is little improvement to be made. The latter might be local minima, where a different discrete path admits a better trajectory.
These improvements require on average an additional 4.08 seconds on top of the 17.28 seconds to generate the original solutions for the Euler angles. Of this time, the solver only was running for 0.62 seconds. The remaining 3.46 seconds were set-up time to re-use the vertices and Bézier curves from the original solution; this time could be further optimized. Comparatively, the time to generate GCS paths when planning with axis-angles is 41.10 seconds. At a lower resolution quaternions take 16.73 seconds, but for higher resolutions, their solve time is on the order of minutes. Our method offers a good payoff between finding shorter paths with Euler angles while still being faster than the more accurate axis-angle and quaternion representations. Euler angles are also currently the only parametrization compatible with the IRIS-NP algorithm [29] for growing collision-free regions in the presence of obstacles.
@@ IV-C Rational Parametrizations of Robot Kinematics
We have two experimental settings that use the rational kinematics parametrization. One is a 3 degree-of-freedom iiwa (four of the joints are locked) that moves within a vertical 2D plane. The other is a 7 degree-of-freedom iiwa mounted on a table, as shown in Figure 1. The nominal position (i.e. point of projection) for both iiwas is when the arms stand straight up with all joint angles at 0. All the regions in the 3DoF case are certified to be completely collision-free using the Certified IRIS algorithm [5]. All the trajectories in the 7DoF setting can be certified using [18].
For the 3DoF planar iiwa, visually the paths become less biased towards the point of projection. For example, in Figure 6, the original paths in red spike towards the nominal position, but after the PGD refinement, the end effector has less extraneous motion. Quantitatively, when measured across 100 randomly chosen start and goal points navigating the shelves, most paths do not have a lot of improvement to be made. On average, the paths get about 0.2% shorter and most terminate within 7 iterations and 0.22 seconds. For the specific path in Figure 6, the numerical improvement is a 1.2% shorter path length. Weaker numerical results for this case is expected as the distortion of configuration space is most intense near the boundaries of the configuration space (i.e. near the joint limits), so the average path does not have a lot of room for improvement.
@@
Figure 6: The 3DoF iiwa (right) shows a decrease in skew towards the nominal position as compared to the original GCS solution (in red). The 7DoF iiwa (left) shows an improvement in path length before and after the post processing for a random selection of start and goal points.
@@ V Discussion
Our method is particularly effective when accounting for the distortion from nonlinear parametrizations of planning spaces. In constrained bimanual motion planning, our post-processing step produces paths that are more balanced between the arms, 20% shorter on average, and 31.02% faster after being time-parametrized. For Euler angles, the paths are 40% shorter on average. In addition to undistorting paths from parametrized spaces, the approach enables the optimization of a path based on general nonconvex objectives. Optimizing over curvature, we find paths with greater radii of curvature, facilitating quicker traversal.
For rational kinematic parametrizations, the lack of significant improvement in path length in the average case suggests that the distortion from the stereographic projection is not usually a major problem. Thus, the benefits of planning in this parametrization of configuration space (enabling rigorous certification) plausibly outweigh the minor costs. In the worst case, our method can produce strong improvements, and when there is not much space for improvement, the solver terminates quickly
An obvious limitation of the proposed methods is the added computation time. We find the time taken by the solver quite satisfactory given that we used a Python based custom PGD. The use of commercial solvers, code improvements, and a compiled language could make a mature implementation much faster. In cases where a solution requires strong guarantees and high quality trajectories (such as in path planning and control of surgery robots), a post-processing step will certainly be worth the additional runtime.
Future work can include larger scale parallelization, especially during the rounding stage. Just as cuRobo has shown incredible results by solving many nonconvex trajectory optimization problems in parallel, our post-processing stage could be integrated with a parallel rounding scheme to get better solutions more quickly. This step could also be used in an Anytime Motion Planning framework [30], where the later parts of a trajectory are refined as the earlier parts are traversed. Another possibility is using the nonconvex objectives with incremental search methods such as GCS* [31].
Another line of work might focus on designing better convex surrogates. These convex surrogates still play an important role during the convex relaxation and initialization stages for the nonconvex optimization. When there are clear deficiencies in the convex surrogate (such as in the original constrained bimanual case), one can try to hand-design a better surrogate. When it is unclear how to make a better surrogate, one could potentially generate them automatically using learning-based approaches.

### v2

@@ III-D2 Optimal Step Sizes
Our objectives are too complex to easily identify the Lipschitz constant and theoretically find a good step size. Classical PGD would then require manually tuning step size, so we use the backtracking line search PGD [27], which searches for an optimal step size. It repeatedly halves an upper bound on the step size till the Armijo condition of sufficient decrease is met. This keeps the solver from overshooting minima while converging fast.
@@ III-D3 Gradient Precompilation
Initially, gradient computations were most of the runtime. Precompiling gradients with JAX [28] moves this time cost offline to speed up the PGD iterations. Compiling gradients for each vertex individually also allows us to re-use computations for start goal pair.
@@ III-E More general nonconvex objectives: Curvature
should help TOPP-RA produce better trajectories, as high curvature paths contain tight turns, that require a slower traversal to stay within acceleration limits. Although such paths might be longer than those produced by a pure shortest-path trajectory, they can be traversed more quickly.
We define this objective too over sampled points along the path defined by $x_{v}$. Given sampled points, we calculate the curvature of each point and then apply the RealSoftMax (a smooth maximum function) to approximate the maximum curvature of our paths. We expect paths under this optimization will have higher path length but lower duration when time-parametrized by TOPP-RA.
@@ IV Experiments
Figure 3: Optimizing jointly for curvature and distance yields quicker trajectories but longer distances–the curvature-regularized path is farther from the shelf.
@@ IV-A Constrained Bimanual Motion Planning
In this experiment, two iiwas navigate a shelf while keeping the transform between end effectors constant, as if they were jointly carrying an object as shown in Figure 1. We evaluate the PGD solver on the key start and goal pairs from [3] both on hardware and in simulation. The comparison of these benchmark paths before and after optimizing for the nonconvex objective is presented in Table I. The “GCS” column indicates using the convex $\mathbb{R}^{8}$ objective, the “Distance” column indicates using the nonconvex $\mathbb{R}^{14}$ objective, and the “Curvature + Distance” column indicates using a linear combination of the nonconvex $\mathbb{R}^{14}$ objective and the nonconvex curvature cost with a ratio of 8 to 0.01 respectively. This ratio is a hand tuned parameter to compensate for the path distance objective being on the order of 100 times greater than the path curvature objective. While the $\mathbb{R}^{14}$ objective results in the shortest paths, regularizing for lower curvature lengthens paths, but shortens traversal times after TOPP-RA’s re-timing. Visually, minimizing this joint objective leads to more rounded paths, as shown in Figure 3.
To quantify the difference in distance traveled between the arms, we define the imbalance of a trajectory as $(d_{s}-d_{c})/(d_{s}+d_{c})$, where $d_{c}$ is the distance traveled by the controlled arm and $d_{s}$ is the distance traveled by the subordinate arm. When both arms travel comparable distances, the imbalance distribution centers around 0. When one arm travels much longer distributions than the other, the imbalance metric approaches ±$\pm 1$ in magnitude. Table I shows that this imbalance metric approaches 0 after post-processing under the $\mathbb{R}^{14}$ objective. In Figure 4, we see paths favour the leading arm less. The imbalance for jointly optimizing curvature and distance is higher than optimizing just the distance indicating that smoother paths are more imbalanced. This asymmetry likely comes from the same-handedness of the iiwas.
For a more comprehensive analysis, we randomly sample 100 start and end points from the valid and reachable configuration space. Paths generated are on average 20.60% shorter in the $\mathbb{R}^{14}$ configuration space after applying our post-processing step. These paths take on average 31.02% less time to navigate. The imbalance shifts towards 0, indicating that the paths for the subordinate arm are more comparable to the leading arm after the nonconvex optimization. These improvements took an average of 0.0554 seconds of compute (approximately 13.7 iterations) in addition to the 2.133 seconds that the surrogate convex optimization takes.
TABLE I: Optimizing over the nonconvex cost improves metrics for the three benchmark trajectories.
@@
Imbalance 0.190 0.084 0.122
Figure 4: Paths become more centered as the nonconvex objective accounts for the distance traveled by both arms. The original convex objective just accounts for the controlled arm.
@@ IV-B Planning over $\operatorname{SO}(3)$
For our last experiment, we plan over random start and goal 3D rotations independent of a simulation or hardware set-up. To cover SO(3), we set up the same charts and convex regions as in [4] for the Euler angles, quaternions, and axis-angle parametrizations. The latter two use piecewise-linear approximations of the original $\operatorname{SO}(3)$ space and act as baselines. We run PGD on the Euler angles setting only. The Euler angles GCS graph is fully connected and optimizes Euclidean distance within each set, so the shortest path between any two points will be a linear path, regardless of the order of our Bézier curves. For time-efficiency, we generate order one GCS solutions and initialize the PGD solver with the control points evenly spaced along each line segment.
@@
Given there are no obstacles, SLERP gives the shortest path between any two orientations. We use it as the closed-form ground truth distance. Figure 5 shows the distribution of relative error in path length for the three representations of $\operatorname{SO}(3)$ when planning across 125 random start and goal pairs, along with the PGD post-processing on the Euler angles paths. The distribution of error for Euler angles shifts significantly closer to 0 after running PGD. The relative error decreases by 42.5% on average. This improvement has a bimodal distribution: for some paths the PGD greatly improves the solution, but for others, there is little improvement to be made. The latter might be local minima, where the global optimal lies on a different discrete path.
These improvements on average take 4.08 seconds in addition to the 17.28 seconds taken to generate the original solutions for Euler angles. Of this time, the solver only ran for 0.62 seconds. The remaining 3.46 seconds were spent re-using vertices and Bézier curves from the original solution and could be further optimized. Comparatively, planning with axis-angles takes 41.10 seconds. At a lower resolution quaternions take 16.73 seconds, but for higher resolutions, their solve time is on the order of minutes. Our method offers a way to generate more accurate paths using Euler angles while still being faster than the more accurate axis-angle and quaternion representations. Moreover, of the three, only Euler angles allow for using IRIS-NP [29] to grow collision-free regions in the presence of obstacles.
@@ IV-C Rational Parametrizations of Robot Kinematics
We have two experimental settings in simulation that use the rational kinematics parametrization. One is a 3 degree-of-freedom iiwa (four of the joints are locked) that moves within a vertical 2D plane. The other is a 7 degree-of-freedom iiwa mounted on a table, as shown in Figure 1. The nominal position (i.e. point of projection) for both iiwas is when the arms stand straight up with all joint angles at 0. All the regions in the 3DoF case are certified to be completely collision-free using the Certified IRIS algorithm [5]. All the trajectories in the 7DoF setting can be certified using [6].
For the 3DoF planar iiwa, qualitatively the paths become less biased towards the point of projection: in Figure 6, the the PGD refinement reduces the extraneous spike towards the nominal pose. Quantitatively, most paths show little improvement across 100 random start and goal points among the shelves. On average, the paths get 0.2% shorter and most terminate within 7 iterations and 0.22 seconds. The example in Figure 6 shows a 1.2% improvement in path length. Weaker numerical results are expected as the configuration space distorts most intensely near the joint limits, so the average case does not have much room for improvement.
@@
Figure 6: The 3DoF iiwa (right) skews towards the nominal position in the original GCS solution (in red). The 7DoF iiwa (left) shows improvement in path before and after the post-processing for a random start and goal configuration.
@@ V Discussion
Our method is particularly effective when accounting for the distortion from nonlinear parametrizations of planning spaces. In constrained bimanual motion planning, our post-processing step produces paths that are more balanced between the arms, 20% shorter on average, and 31.02% faster after being time-parametrized. For Euler angles, the paths are 40% shorter on average. Beyond undistorting paths, the approach enables optimizing general nonconvex objectives such as curvature. For the bimanual setting, we find paths with greater curvature radii and quicker traversal. The lack of significant change in path length in the average for the rational kinematic case suggests that the distortion from the stereographic projection is not usually significant. Thus, planning in this parametrization of configuration space and enabling rigorous certification plausibly outweighs the minor cost increase. Even then, our method produces strong improvements in the worst case, and in the average case with little room for change, the solver terminates quickly.
[…]
An obvious limitation of the proposed method is added computation time. We use a Python based custom PGD; Commercial solvers, compiled languages, and performance optimization will speed up a mature implementation. This post-processing step will certainly be worth the additional runtime in cases like surgical robots that require strong guarantees and high quality. We have focused our numerical results on sparse environments. Our method works in dense clutter (see Figure 7) given IRIS regions generated using new methods that better scale with environment complexity [30]. But the regions are smaller, leaving less room for improvement. In many contexts, robots move through areas of dense and sparse clutter, and our method can improve segments in the sparser regions, without adding collisions or worsening the trajectory in the densely-cluttered areas.
Future work could include larger scale parallelization, especially if we integrate our post-processing step into the rounding stage. cuRobo [17] has shown incredible results by solving many nonconvex trajectory optimization problems in parallel. This step could also be used in an Anytime Motion Planning framework [31] where the later parts of a trajectory are refined as the earlier parts are traversed. Another possibility is using the nonconvex objectives with incremental search methods such as GCS* [32] and Multi Query Shortest Path Problem in GCS [33]. Lastly, we work on designing better convex surrogates which still play an important role during the convex relaxation and initialization stages. Under clearly deficient convex surrogates (such as in the original constrained bimanual case), one can try to hand-design a better surrogate or potentially generate them automatically using learning-based approaches.